### v1

# A Longitudinal, Multinational, and Multilingual Corpus of News Coverage of the Russo-Ukrainian War

Abstract
We introduce dnipro, a novel longitudinal corpus of 246K news articles documenting the Russo-Ukrainian war from Feb 2022 to Aug 2024, spanning eleven media outlets across five nation states (Russia, Ukraine, U.S., U.K., and China) and three languages (English, Russian, and Mandarin Chinese). This multilingual resource features consistent and comprehensive metadata, and multiple types of annotation with rigorous human evaluations for downstream tasks relevant to systematic transnational analyses of contentious wartime discourse. dnipro’s distinctive value lies in its inclusion of competing geopolitical perspectives, making it uniquely suited for studying narrative divergence, media framing, and information warfare. To demonstrate its utility, we include use case experiments using stance detection, sentiment analysis, topical framing, and contradiction analysis of major conflict events within the larger war. Our explorations reveal how outlets construct competing realities, with coverage exhibiting polarized interpretations that reflect geopolitical interests. Beyond supporting computational journalism research, dnipro provides a foundational resource for understanding how conflicting narratives emerge and evolve across global information ecosystems.
A Longitudinal, Multinational, and Multilingual Corpus of News Coverage of the Russo-Ukrainian War
Dikshya Mohanty, Taisiia Sabadyn, Jelwin Rodrigues, Chenlu Wang, Abhishek Kalugade, Ritwik Banerjee Stony Brook University Correspondence: {dimohanty, jelrodrigues, rbanerjee}@cs.stonybrook.edu
1 Introduction
The Russo-Ukrainian war, which escalated on February 24, 2022 with the declaration of a "special military operation" by the Russian government, [^1] Within 48 hours, media in Russia were banned from describing the developments as “war” AFP (2022). has significantly impacted the global geopolitical landscape Derviş (2023); Silvertsen (2024). The war has redrawn traditional alliances, tested the resolve of international institutions, and created new geopolitical fault lines. Despite initial expectations of a swift Russian victory, Ukraine’s resistance has resulted in a protracted war with substantial loss of life and human capital Watling et al. (2023); Pancevski (2024). This conflict has also been characterized by a quieter battle for narrative control, with news coverage from several nations playing a pivotal role in shaping global public opinion and influencing the war’s trajectory Bradshaw et al. (2024). These sources reflect their national priorities, strategic perspectives, and the perceptions they intend to create among information consumers Johais and Meis (2024). Yet, despite the centrality of media and narrative framing, there remains a lack of comprehensive datasets for longitudinal analysis of how nations have framed the war for their domestic and international audiences.
We introduce a collection of Diverse Narratives and International Perspectives on the Russo-Ukrainian Offensive (dnipro), a comprehensive longitudinal corpus of news coverage from February 2022 to August 2024222An expanded version is being developed, which will include news coverage up to February 2025 and provide a longitudinal 3-year corpus of this protracted war.—the longest time span of any multinational or multilingual news corpus on the Russia-Ukraine war. The corpus encompasses eleven influential media outlets and news agencies aligned with five primary geopolitical actors with strategic interests in the conflict. Following methodological approaches established by Nygren et al. (2016) and Roman et al. (2017), the sources include media from direct participants—Russia and Ukraine—and influential global powers with economic, diplomatic, or military involvement: USA, UK, and China. The inclusion of state-affiliated media from Russia and Ukraine capture opposing conflict narratives, while those from China, the USA, and the UK represent the narratives projected by nations with significant stakes in the conflict’s outcome, and significant ability to influence that outcome Zhang et al. (2016); Thussu et al. (2018).
dnipro offers rich annotations for entity-level stance, sentiment, and topical framing, enabling event tracking and other longitudinal analyses across geopolitical perspectives. The dataset does not attempt to verify the accuracy of any claims; rather, it provides a resource for studying how wartime narratives are constructed and disseminated by major geopolitical actors. [^3] This aligns with established methodologies in discourse analysis Entman (1993); Fairclough (2013) examining how language and framing construct social reality. Further, Gamson and Modigliani (1989) showed how this view offers a powerful tool to understand the competition that shapes public understanding of contentious events. The corpus is uniquely suited to such studies, as it comprises 246,229 articles across 11 media sources and 3 languages, spanning 5 diverse nations, over a period of 31 months. Our key contributions include:
  • (1)
    a longitudinal multinational and multilingual corpus of news coverage spanning the first 30 months of the Russia-Ukraine war;
  • (2)
    annotations supporting comparative analyses of named entities, stance, and framing; and
  • (3)
    the inclusion of diverse media sources enabling cross-cultural examination of geopolitical narratives.
Next, we place dnipro within the landscape of existing conflict-related corpora and highlight its distinct contributions to the field (§ 2), and describe how our design and construction choices (§ 3) make it a powerful resource for large-scale analyses of wartime narratives. We employ dnipro in several applications, using state-of-the-art language understanding systems as well as additional annotations (§ 4 and § 5), demonstrating its utility.
2 Related Work
Research on wartime media coverage relies on datasets capturing narratives across national and institutional contexts, where competing claims and propaganda techniques pose unique challenges. Such resources are vital for training models to analyze geopolitical discourse and evaluating their ability to identify narrative strategies. As a reflection of diverse goals, existing resources vary in scope, modality, and annotations.
Twitter & Reddit-based datasets are among the largest. Chen and Ferrara (2023) aggregate over 570 million tweets in English, Russian, and Ukrainian to study information propagation and influence operations, while Zhu et al. (2022) provide 8 million posts correlating post categories with pro-Ukrainian stance. Shevtsov et al. (2022) introduce 57.3 million English tweets annotated for sentiment and hate speech, and the Invasion@Ukraine corpus provides 8.7 million English tweets for content moderation and conspiracy narrative analysis Susanne Pohl et al. (2023). The VoynaSlov collection Park et al. (2022) bears similarities to our work through sentence-level media framing labels, though they focus solely on Russian media.
Alyukov et al. (2023) introduce one of the first cross-platform social media datasets, integrating 1.7 million posts to enable propaganda and public opinion analyses. Telegram-specific corpora include 4.1 million posts with pro/anti-Kremlin channel tags Bawa et al. (2025) and 2.3 million posts with binary tags for human rights violations Nemkova et al. (2023).
Hakimov and Cheema (2024) introduce a corpus of 1.5 million tweets in 60 languages with automated sentiment, stance, and named entity annotations, offering the widest multilingual social media dataset. The RUWA corpus Khairova et al. (2024) provides 16,500 news articles using semantic similarity for misinformation detection, but excludes the overarching war theme and curates only specific event-based articles. dnipro supersedes it by orders of magnitude in coverage, topics, multilinguality, temporality, and size. Ibrahim et al. (2025) offer longitudinal media analysis of two years of news data, but from nations with auxiliary involvement.
Annotation practices differ widely: manual schemata for framing Park et al. (2022), channel stance Bawa et al. (2025), and human rights violations Nemkova et al. (2023); automated annotations for large-scale labeling Susanne Pohl et al. (2023); Hakimov and Cheema (2024). Most resources rely on social media identifiers that degrade quickly, impeding reproduction. Few provide longitudinally contiguous, multilingual, news-based content annotated for multiple tasks.
dnipro extends existing corpora by focusing on transnational and multilingual news coverage – underrepresented in prior social media-centered work – enabling comparative analyses of state-aligned narratives across geopolitical actors. It reflects news coverage as disseminated, without alteration, serving as a neutral resource for empirical studies. Our design choices spanning source selection, temporal coverage, and structured annotations make dnipro a powerful resource for large-scale automated analysis of wartime narratives (§ 5).
3 The dnipro Corpus
Our dataset is constructed through a systematic multi-stage pipeline targeting news outlets and state-affiliated agencies selected on the basis of three criteria: (1) global influence and prominence within national media ecosystems, (2) sustained coverage of the Russo-Ukrainian war, and (3) discernible editorial positioning ranging from independent journalism to state-aligned messaging. To the extent possible, access to full article archives was obtained through public RSS feeds, institutional subscriptions, or source-specific data collection pipelines. However, not all media outlets provided consistent or unrestricted access to their archives, and some selections reflect practical constraints in addition to topical relevance.
3.1 Data Acquisition
dnipro is a corpus spanning 31 months of daily data collection, from February 1, 2022 to August 31, 2024. To identify articles with substantive coverage of the Russo-Ukrainian conflict, we employ a tiered strategy. For outlets with dedicated sections or labels related to the war, we collect all articles published under those categories. Where available, we also extract articles tagged with conflict-specific metadata (e.g., “Ukraine conflict”, “special military operation”, “Russia-Ukraine war”). For sources lacking such taxonomies, we apply keyword-based filtering to isolate relevant content. However, some sources impose significant restrictions on retrospective API-based keyword searches. In these cases, we obtain the older articles directly through web browsers, operating within the scope of respective terms of service and copyright restrictions. To ensure quality control, we manually validate stratified random samples across sources, languages, and time periods throughout the acquisition process.
Total no. of articles 246,229
Title length (avg.) 11.7 words
Article length (avg.) 332.5 words
No. articles/day (avg.) 261.1
No. of sources 11
Languages covered N/A
Temporal span Feb 1, 2022 - Aug 31, 2024
No. of non-English articles 61,637 (25.03%)
Geopolitical entities Russia, Ukraine, China, USA, UK
After translation (of non-English sources to English).
Table 1: Key descriptive statistics of dnipro.
English-language “Western” media from the US and the United Kingdom, namely The Financial Times (FT) and Cable News Network (CNN), were included as part of the Western bloc, reflecting the perspectives of countries with significant political, economic, and military involvement in the conflict. These outlets serve domestic and international audiences and are central to understanding how Western-aligned narratives are constructed and disseminated. No additional filtering of articles is done by search queries as these media houses offer dedicated sections for content related to the Russo-Ukrainian war. We collect a total of 5,760 and 4,218 articles from CNN and FT, respectively.
To ensure adequate representation of Chinese state-aligned perspectives, we collect articles from three leading outlets: Xinhua, Global Times, and China Daily. Xinhua publishes exclusively in Mandarin, whereas Global Times and China Daily provide parallel English-language coverage. Articles are collected from Xinhua using a bilingual keyword list, designed to minimize the loss of coverage while retaining topical precision. Mandarin Chinese queries comprised “乌克兰” (Ukraine), “俄罗斯” (Russia) and “俄乌” (Russia and Ukraine); English queries included “Russia–Ukraine”, “Russia”, “Ukraine”, and “Ukrainian”. These queries yield in 11,152 and 10,206 articles in Mandarin Chinese and English, respectively.
[…]
The Ukrainian perspective is represented by articles from Censor.net, The Kyiv Independent (KI), and European Pravda (EP), all of which provide English-language versions. While KI and EP shape Ukraine’s global image, Censor influences the domestic audience. Their reporting styles vary—from Censor’s rapid updates to EP’s in-depth opinions and interviews. As sources targeting the global Anglosphere, KI and EP offer sections dedicated to the Russo-Ukrainian War: “War” and “War with Russia”, respectively. Censor’s news reporting approach largely covers the war, evidenced by the majority fraction of Ukrainian war-related tags in their “News” section. As such, no keyword search queries are applied to any of the Ukrainian sources. The collection from these courses amount to 19,706 articles from Kyiv Independent, 2,212 from European Pravda, and 31,748 from Censor.
The Russian media sources in our dataset comprise English articles from TASS and Sputnik, and Russian-language articles from Izvestia. No keyword search queries are required, as all articles are collected from the dedicated section “Special Operation of Russia in Ukraine”. We collect 60,002 and 53,139 articles from TASS and Sputnik respectively, and 50,485 articles from Izvestia.
article_id A unique UUID v4 identifier for the record.
url The persistent URL from which the article was sourced.
publication_date The article’s publication date in ISO 8601 format (yyyy-mm-dd).
source_name The normalized name of the news outlet (e.g., Xinhua).
source_country The country of origin for the news outlet, in ISO 3166-1 alpha-3.
language The article language, in ISO 639-1.
title The article’s title (i.e., headline).
article_length The number of words in the cleaned article text (post processing).
collection_date The date (in ISO 8601 format) when the URL was accessed.
Table 2: Core metadata fields for each article.
3.2 Metadata Schema and Statistics
The core of dnipro is a collection of rich, consistent metadata records for 246,229 news articles. Its key descriptive statistics are shown in Table 1. Each record is described by a unified schema (described in Table 2) designed to facilitate large-scale longitudinal and comparative analyses.
The overall composition of the natural language data—both in terms of the number of news articles and their sizes—across individual news sources and aggregated by geopolitical entities, is presented in Figure 1.
3.3 Data Records
dnipro is publicly available on Zenodo under a CC-BY-NC 4.0 license. [^4] The data, along with all processing code and annotation prompts, will be released upon acceptance. It is structured for ease of use and distributed as a single, compressed tar.bz2 archive containing a primary metadata file, annotation files, and a documentation file. [^5] We do not directly include the texts due to copyright restrictions. The provided URLs serve as persistent identifiers enabling reproduction of dnipro, like most social media datasets. If an article becomes unavailable, the collection_date and url enable retrieval via the Internet Archive’s WayBack Machine.
  1. 1.
    metadata.parquet contains the complete set of 246,229 metadata records as described in § 3.2. The Parquet format was chosen for its efficient compression and columnar storage, which enables fast querying and analysis using popular data science tools.
  2. 2.
    named-entities.parquet contains the named entity annotations for the entire corpus. Each record includes an article_id (see Table 2) and the list of named entities in that article.
  3. 3.
    sentiments.parquet provides sentiment scores for select entities and events across multiple news sources. For politically important and frequently occurring entities (e.g., Ukraine, Russia, Zelensky, Kherson), these scores are aggregated monthly. For entities tied to key events within this war (e.g., Bucha, Kursk, Zaporizhzhia), they are aggregated daily.
  4. 4.
    topical-frames.parquet provides one row per article_id with its corresponding topical frames anchored in the nine-category schema described in Table 3.
  5. 5.
    README.md is a comprehensive description of the dataset, the schemata, the license terms, and a code of conduct for ethical use.
We do not specify train/test/dev splits for dnipro. The corpus spans time, languages, and multiple sources; and supports diverse study designs (e.g., temporal evolution or development of an entity/topic/event; media source comparisons; transnational or cross-lingual reporting of events; or other types of aggregation). A preset split would constrain the diversity of possible applications. We encourage task-appropriate partitions instead.
4 Data Annotations
We now describe the enrichment of raw article metadata with annotations. Since copyright restrictions prevent the direct release of texts, we generate content-derived labels that form the dataset’s core analytical value. The process involves:
  1. 1.
    metadata and text normalization;
  2. 2.
    translation of non-English text to English;
  3. 3.
    labeling named entities;
  4. 4.
    sentiment analysis;
  5. 5.
    stance analysis;
  6. 6.
    topical framing; and
  7. 7.
    rigorous quality control with human evaluation.
4.1 Text Processing for Annotation
To prepare the corpus for annotation and analyses, we apply standardized processing steps to all data, across sources and languages. This ensures consistency across media sources with varying conventions in structure, formatting, and metadata.
Publication dates are parsed and normalized to ISO 8601 standard (yyyy-mm-dd), and multiple metadata fields—e.g., “publication date”, “published on”—are mapped to metadata fields in our unified schema (see Table 2). Text content is stripped of non-informative boilerplate elements like location or date prefixes commonly found in news articles (e.g., “KYIV, March 3 –”), inline footnotes, bylines, and repeated source disclaimers. Duplicate articles are removed, to avoid redundancy in print syndication and multilingual reprints. To support downstream linguistic analyses, the text is then converted to UTF-8 encoding, segmented into paragraphs, and split into sentences.
Frame Definition
economy Macroeconomic issues and markets: sanctions/trade, inflation, growth, jobs, taxes, energy/commodity prices.
identity Group identity and belonging: nationality, ethnicity, religion, language, or gender; discrimination or identity politics.
morality Explicit ethical/value framing: moral duty, fairness, dignity, humanitarian arguments.
legal Courts and legal processes: rulings, lawsuits, prosecutions, designations, warrants, legal rights or entitlements.
policy Concrete government outlook or outreach, rules, bills, decrees, program design and details (what a policy does).
politics elections, party/elite competition, horse-race strategy, reshuffles, coalition dynamics.
public opinion Polls, surveys, attitudes and protests as expressions of public sentiment.
security War/defense, violence/crime, policing, border enforcement, terrorism, cyberattacks.
welfare Benefits and human services: healthcare, pensions, housing aid, unemployment, assistance to refugees or internally displaced people.
Table 3: Topical frame definitions in dnipro: a significant abridgment of the Refugees and Migration Framing Vocabulary Yu and Fliethmann (2022).
4.2 Machine Translation
A quarter of dnipro consists of Mandarin Chinese and Russian texts, (zh and ru, respectively, in ISO 639-1). We translate these into English (en) to facilitate downstream analyses. We choose three models based on recent comparative studies on neural machine translation by Smirnov et al. (2022) and Alemayehu et al. (2024): OPUS-MT Tiedemann et al. (2023), M2M100 Fan et al. (2020), and NLLB-200-Distilled Team (2022). Since there were no readily available explicit benchmarks for zh-to-en or ru-to-en translations, we select a random set of ten articles from Xinhua, evaluated by two native speakers of Mandarin Chinese for semantic fidelity and fluency. The evaluators worked independently to score the output of these models on 1-5 Likert scale for fidelity and fluency, and M2M100-418M received the highest scores from both evaluators on both criteria. So, we employ this model for all zh-to-en translations. Russian articles were similarly translated using the ru-to-en path of M2M100-418M.
4.3 Annotation Methodology
To enable fine-grained analysis of key actors and locations, we perform Named Entity Recognition (NER) across the entire corpus (after translation to English). This annotation is carried out using the spaCy library Honnibal et al. (2020) and its pretrained model en_core_web_sm, identifying entities and their types according to the OntoNotes 5.0 taxonomy Weischedel et al. (2013). This process results in the identification of 85,155 distinct named entity mentions, and provides a foundational layer for tasks such as tracking the prominence of specific actors over time, temporal studies of media focus across entities and events, and entity-specific sentiment and stance analyses.
Event (Date range) # articles
Bucha Massacre (Apr 3–9, 2022) 273
Zaporizhzhia Nuclear Crisis (Mar 3–9, 2022) 219
Mariupol Hospital Attack (Mar 9–15, 2022) 138
Kursk Offensive (Aug 6–12, 2024) 180
Table 4: The four conflict events analyzed in our use case experiments, selected based on global coverage and the potential to generate narrative divergence.
To capture the sentiment of some of the most prominent named entities, we provide sentence-level sentiment annotation using a pretrained model: sentiment-roberta-large-english Hartmann et al. (2023). For each article, sentences containing such entities (e.g., Kursk, Bucha, Kherson, Zelensky, Ukraine, Russia) are identified using spaCy, and passed to this classifier. The model outputs are then mapped to continuous sentiment scores, which are then aggregated on a daily and monthly basis at both the source and country level using source–country relations. These annotations thus provide time series data revealing sentiment trends for each entity.
The dataset includes stance annotations obtained by few-shot instruction prompting with Phi-4 14B model  Microsoft (2024), identifying the stance towards a specified target. The input comprises a sentence and the immediate context in its news article. As part of the input, we provide a task instruction with a few examples, along with (i) a passage for context resolution, (ii) a focal sentence, and (iii) the target. The model is prompted to output a compact, formatted response:
Target: <the identified target>
Stance: favor|against|none
Rationale: <a short justification grounded in the given text>
For topical framing, we again use few-shot prompting with Phi-4 14B model, where the input is an instruction describing the task, along with label definitions and 1-2 examples per label. Then, for a given instance – an article’s title, subtitle (if present) and content – we prompt the model to generate a json object containing the list of applicable topical frames, and associated confidence scores.
4.4 Human Evaluation
To evaluate the reliability of our automated annotations, we conduct human evaluations of the stance and topical frame annotations.
Frame PABAK Confidence
Human/Human Human/LLM Human LLM
Economy 0.70 (84.8%) 0.736 0.84 0.93
Identity 0.74 (86.8%) 0.783 0.79 0.92
Legal 0.85 (92.4%) 0.783 0.74 0.91
Morality 0.75 (87.3%) 0.750 0.79 0.91
Policy 0.55 (77.7%) 0.526 0.78 0.89
Politics 0.48 (74.1%) 0.425 0.79 0.92
Public Opinion 0.87 (93.4%) 0.865 0.80 0.89
Security 0.69 (84.3%) 0.587 0.85 0.93
Welfare 0.81 (90.4%) 0.783 0.80 0.87
Average 0.71 (85.7%) 0.693 0.798 0.909
Table 5: Inter-annotator agreement and confidence by topical frame. Pabak columns report human inter-annotator agreement (percent agreement in parentheses) and human–model agreement. Confidence columns report average confidence scores from human annotators and LLM (Phi-4).
Topical Frames:
Two expert annotators independently label a stratified random sample of 200 articles, providing nine binary labels with corresponding [0,1] confidence scores for each article. Significant class imbalance is a common challenge in multi-label annotation, and the inter-annotator agreement must consider the high base rate agreement where annotators agree on a label being absent from an article. The average agreement on label absence, over all nine topical frames, accounts for 65.82% of our evaluation sample, with the positive class (for any label) representing a minority of instances. Due to this prevalence bias, traditional measures like Cohen’s or Fleiss’ kappa are overly pessimistic reliability measures for our data. For a more accurate assessment that accounts for this imbalance, we report the prevalence-adjusted bias-adjusted kappapabak Byrt et al. (1993) – which is especially crucial for evaluation of consensus on the rarer labels.
The inter-annotator agreement results, measured by pabak, are presented in Table 5 and demonstrate substantial agreement on average (0.71, with 85.7% agreement). Three topical frame categories — Welfare, Public Opinion, and Legal — reveal near-perfect agreement (>> 0.8). Overall, seven of the nine labels achieve pabak scores $\geq$ 0.7, indicating a consistent and shared understanding of the annotation framework. The two remaining classes, Policy and Politics, have moderate agreement, as they represent more abstract and often overlapping conceptual domains. We also examine the correlation between the average confidence scores reported by the human experts and those reported by Phi-4. The framing categories are complex and sometimes overlapping, resulting in a weak positive correlation $\rho$ = 0.37.
Implications for dataset use. The strong agreement validates the overall reliability of our annotation schema. However, we recommend that users consider the tiered nature of this reliability when designing their experiments:
  • The seven frames with pabak > 0.69 are suitable for fine-grained analysis, including temporal tracking and cross-source comparison.
  • The Policy and Politics labels yield lower agreement due to their inherent conceptual overlap. We recommend these labels be used for high-level topic modeling or as supplementary features in larger models, with the understanding that they represent broader and possibly intersecting, categories of discourse.
  • The human-annotated sample of 200 documents, while modest in size, provides a high-quality validated benchmark. It was strategically selected through stratified sampling to ensure longitudinal representation across sources and languages, making it suitable for evaluating the performance of topical frame classification systems on real-world data on contentious wartime narratives.
Annotator-pair Agreement (%) Cohen’s $\kappa$
Human annotators 86.00% 0.770
Annotator1 and phi-4 72.68% 0.618
Annotator2 and phi-4 70.00% 0.580
$\bullet\;\;$ Mean agreement: 76.23%
$\bullet\;\;$ Fleiss’ $\kappa$ across all three annotators: 0.643
$\bullet\;\;$ Mean pairwise Cohen’s $\kappa$: 0.656 (range: 0.58-0.77)
Table 6: Inter-annotator agreements for stance labels, measured by percentage of agreement, pairwise Cohen’s $\kappa$, and Fleiss’ $\kappa$.
Stance Annotations:
The stance annotation task identifies whether a sentence (with context) expressed a stance (favor or against) towards any entity, or if no stance was present (none). Two human annotators with native proficiency in English annotated a sample of 150 instances. We evaluate their inter-annotator agreement, and then also with phi-4 as the LLM annotator. Table 6 presents the pairwise and overall inter-annotator agreements, which indicate strong human inter-annotator agreement ($\kappa$ = 0.77). The LLM annotator achieved moderate agreement with both ($\kappa$ = 0.62 and 0.58), showing consistent but less-than-human reliability. Fleiss’ $\kappa$ = 0.643, however, suggests that the stance annotations are sufficiently reliable for large-scale analysis, despite the gap between human and automated annotation quality for this task.
Figure 2: Topical framing by source, with % share of articles for each label (confidence >> 0.5). Source abbr.: Financial Times (FT), Kyiv Post (KP), Sputnik (SP), European Pravda (EP), Global Times (GT), China Daily (CD), Xinhua (XH). Labels: (Sec)urity, (Pol)itics, (Econ)omy, (Welf)are, (Ident)ity, (Moral)ity, and (Pub)lic (Op)inion.
5 Use Case Experiments
To demonstrate the analytical potential of dnipro, we present several studies that leverage its unique multi-perspective structure. These experiments are intended as use case demonstrations of the dataset’s capacity to support diverse research questions in computational journalism, political science, and discourse analysis. Each application highlights specific affordances of the corpus while suggesting broader research directions.
Thematic Analysis of Conflict Events:
Our analyses focus on four major conflict events that received substantial global coverage and generated clear geopolitical divisions, inviting transnational comparisons (see Table 4). For each event, we filter articles using specific location identifiers (e.g., Bucha) to ensure event-specific analysis.
To examine how media outlets prioritize different themes when covering conflict events, we analyze the topical content and emphasis patterns across sources. We identify which concepts, actors, and subtopics receive attention in coverage of major events, revealing how outlets differentially emphasize aspects such as historical context, humanitarian concerns, or institutional responses.
Using GPT-3.5 Ye et al. (2023) with structured prompts, we tag articles from weekly time windows surrounding each major event to identify key concepts, actors, and subtopics. For the events described in Table 4, we obtain 810 articles, with 10.5 such tags (avg.) per article. We aggregate these tags by event and source country to find the distinct thematic emphasis patterns across media outlets.
Russian and Chinese outlets consistently emphasize institutional terminology (special military operation), security concerns (nuclear weapons), and counter-narratives (media manipulation, ceasefire), reflecting strategic adversarial framing. Ukrainian sources predominantly highlight humanitarian dimensions (civilian casualties) with frequent references to specific conflict locations. American and British media balance humanitarian topics with institutional responses (sanctions, international criminal court) and policy discussions (no fly zone).
These patterns demonstrate how dnipro captures differences in thematic prioritization across geopolitical actors even when covering the same underlying events.
Russian Media Ukrainian Media
Russia had testimony from residents of Bucha and Mariupol, who reported brutal crimes, such as torture, committed by Ukrainian nationalists. (TASS) All eyewitnesses noted the systematic nature of the killings in Bucha by the Russian army. (Censor)
Russian forces jointly with border guards repelled these attacks and hit enemy positions near the border and in the Sumy Region. (TASS) Russian forces attacked six border areas and settlements of Sumy Oblast on June 7, firing 14 times and causing at least 54 explosions, the Sumy Oblast Military Administration reported. (Kyiv Independent)
Table 7: Contradictory claims from opposing media outlets, highlighting how conflicting narratives are constructed not through direct refutation, but via divergent attribution, framing, and selection biases.
Topical Framing:
We conduct a corpus-wide analysis using our frame definitions (Table 3). Article-level topical frames provide a comparable map of what each media outlet emphasizes, revealing systematic priorities that contextualize downstream analyses such as stance, sentiment, or discourse. Articles are annotated with phi-4 using a few-shot instruction prompt, producing multi-label assignments along with confidence scores.
Figure 2 shows that Security dominates across all sources, with the highest shares in Ukrainian media, followed by Russian state media and Western sources like CNN. The second-most prevalent category is Politics, elevated in Russian and Chinese sources. It is substantially present in other sources too. Economy is most visible in Chinese media and, naturally, The Financial Times, while Policy appears steadily across sources, with higher emphasis in Chinese and British coverage. Relative to others, Russian media leans more on Legal topics. Notably, discussions of Morality are extremely rare despite the focal topic of this dataset being an extended war, indicating that operational, strategic, and material considerations far outweigh explicit ethical rhetoric in war coverage.
Figure 3: Daily average of stance scores towards Ukraine, grouped by media sources from five nations. The first 15 days of the war (left), contrasted with the last 15 days collected for dnipro (right).
Stance analysis:
Based on our stance annotations in § 4.3, we provide a comparison of the stance towards Ukraine between the first fortnight of the war and the last 15 days of dnipro data. This reveals a striking temporal shift across geopolitical lines. In Figure 3, we see that Chinese media was slightly pro-Ukrainian. By August 2024, however, this shifted to mild opposition. Russian coverage is consistently negative, but there is a drastic negative shift. The US media stance becomes even more supportive, while UK news begins to exhibit greater variability. This temporal comparison demonstrates how dnipro can provide analytical insights into reporting patterns. In this use case, it illustrates the polarization of geopolitical alignments.
Divergent Narratives:
To investigate how media outlets construct and contest narratives around specific flash points in the Russo-Ukrainian war, we analyze claims to identify conflicting representations of key events. Building on prior work in conflict journalism Entman (2008); Thussu et al. (2018), this analysis unearths how opposing perspectives are embedded in individual claims.
Using weekly data subsets for each event, we extract sentence-level claims and categorize them into six types adapted from prior factuality and attribution typologies (Kotonya and Toni, 2020; Spinde et al., 2021):
  • (1)
    direct statements of observable or confirmed events (factual);
  • (2)
    claims attributed to specific actors (attributed);
  • (3)
    explicit denials or rebuttals (disputed);
  • (4)
    indirect accounts from third-party media (reported);
  • (5)
    official statements of intent or directive (command/policy); and
  • (6)
    interpretive or evaluative judgments by experts or institutions (analytical/assessment).
With GPT-3.5 plus human-in-the-loop validation, we label claims via structured prompts, and group semantically related claims through agglomerative clustering on Sentence Transformer representations Reimers and Gurevych (2019).
This approach reveals distinct narrative patterns across geopolitical boundaries: Russian sources might characterize an airstrike as precision targeting of military assets (factual/attributed), while American, British, or Ukrainian outlets describe the same event as civilian infrastructure targeting, often citing eyewitnesses (disputed/reported). Analyzing claim pairs shows entailment occurring primarily within aligned media ecosystems: Ukrainian and Western sources reinforce each other through overlapping factual, reported, or command/policy claims, while Russian media maintains internal consistency through attributed and analytical claims citing official statements. Contradictions predominantly emerge across geopolitical divides as competing interpretations rather than direct disputes. Outlets rarely offer explicit rebuttals, instead employing different topical framing, omission, or counter-attribution strategies.
Table 7 presents contradictory claim pairs from Russian and Ukrainian media. While referring to the same events, the claims construct incompatible interpretations of responsibility and action through attribution and selective emphasis.
6 Conclusion
The dnipro corpus represents a significant advancement in resources for computational analysis of media narratives during international conflicts. By integrating diverse geopolitical perspectives, consistent translations, and structured metadata across 246K articles, this dataset enables systematic study of narrative divergence that previous resources could not support. Our exploratory applications demonstrate dnipro’s utility across multiple analytical dimensions, from stance and sentiment detection to topical framing and contradiction analysis. The findings reveal how global media ecosystems construct parallel realities through attribution choices, selective emphasis, and strategic framing, rarely engaging with competing narratives.
Beyond its utility in computational journalism, dnipro provides a foundation for developing more sophisticated methods to detect cross-cultural narrative divergence, identify emergent propaganda strategies, and understand information warfare dynamics. As the global information landscape grows increasingly complex, such resources become essential for technical advancement as well as a broader understanding of narratives that shape our perception of international conflicts. limitations and ethics don’t count towards the page limit
7 Limitations
The Russo-Ukrainian War has not concluded. The dataset, which currently spans from Feb 2022 to August 2024, may omit important narratives and developments resulting from politically significant events thereafter (e.g., foreign policy changes of a powerful geopolitical entity). For the chosen news sources, a wider variety of publications representing NATO nations would be beneficial. Currently, only two of the thirty-two NATO nations are represented (the United Kingdom and the United States of America). There are other media sources like DW in Germany, which are highly referenced globally, so they are an influential factor. Additionally, representation of more countries with ambiguous involvement in the war would add value to a truly global version of the dataset. Potentially interesting countries would be Hungary and Romania, both of which border Ukraine; Belarus, also bordering Ukraine, but known to be an ally of the Russian state; Turkiye, a powerful regional actor, especially in the Black Sea; and Georgia, a NATO contender that continues to maintain reasonably strong relations with Russia.
Another facet limiting dnipro is the imperfect ground-truth annotations for stance, and other discursive phenomena. While this is not a limitation for qualitative analyses or mixed-method approaches, it does hinder nuanced supervised learning directly from the corpus. Our exploratory analyses, while illustrative, are not exhaustive and serve primarily to demonstrate the dataset’s potential rather than to make definitive empirical claims.
8 Ethical Considerations
dnipro contains potentially sensitive content related to an ongoing conflict with intense geopolitical implications. We acknowledge several potential ethical dimensions:
  • Data collection: All content was gathered from publicly available news sources without breaching access restrictions or terms of service. We do not include any copyright-protected content or any private communications.
  • Bias, representation, and inclusion: Our corpus inherently reflects the biases present in global media coverage. The inclusion of opposing narratives—including state-controlled media—is a deliberate choice, to empower comparative analysis. We strongly caution users against interpreting the mere inclusion of a source as endorsement of its content.
  • Potential misuse: While designed for media analysis, computational linguistics, and conflict studies, this resource could also be used for propaganda analysis. And thus, it could also be misused for propaganda analysis. We strongly urge the community to utilize this corpus for constructive research into understanding narrative differences to bridge societal gaps, rather than weaponize any findings.
  • Lost in translation: Machine translation introduces potential distortions, particularly with politically charged terminology and diplomatic language use. We document the details of our translation process. Further, we retain the original articles for enhanced scrutiny and possible resolution by experts in Russian, Ukrainian, and Mandarin Chinese.
  • Contextualization: We provide robust metadata (publication date, source country, outlet ownership) to ensure articles can be interpreted within their proper contexts rather than in isolation.
By openly addressing these considerations, we aim to enable valuable research on contentious discourse while minimizing potential harms. It is our expectation that through the study of such discourse, researchers will be able to bring empirical enhancements to conflict resolution mechanisms. Applications of such datasets are continuously evolving, as is the dataset itself. As such, we earnestly welcome community feedback on additional safeguards.
Appendix A Appendix
A.1 Prompt Design Details
For our quantitative analysis, we used the mistral-small-latest model via the Mistral API to jointly extract stance targets and classify stances, and employed the siebert/sentiment-roberta-large-english model from Hugging Face to estimate sentiment polarity toward named entities. We used GPT-3.5 (via the OpenAI API) to assist with annotation for three tasks: article-level tag extraction, claim identification, and claim type classification. Prompt configurations varied by task. We employed both zero-shot and few-shot strategies, along with system-level instructions and output formatting constraints, following best practices in prompt engineering for language models Saravia (2022).
A.1.1 Annotation for Stance
For stance detection, we used the mistral-small-latest model via the Mistral API. The task was to classify whether a piece of news content expressed a stance FAVOR, AGAINST, or NONE toward a generated target. We followed a combined target generation at first followed by stance classification.  Akash et al. (2025) We ran these experiments on 501 articles that mentioned the Bucha massacre dating from March to April 2022.
The following zero-shot prompt was used to guide the model:
Prompt:
Listing 1: Zero-shot prompt for joint target generation and stance classification.
Stance classification is the task of determining the expressed or implied opinion, or stance, of a statement toward a specified target. Analyze the following text, generate the target for this text, and determine its stance toward the generated target. A target should be the topic discussed in the text. The target can be a single word or a phrase, but its maximum length must be 4 words.
If the stance is in favor of the generated target, write FAVOR; if it is against the target, write AGAINST; if it is ambiguous, write NONE. The answer must be exactly one of these three words: FAVOR, AGAINST, or NONE.
Output format:
Target: <target>, Stance: <stance>
A.1.2 Annotation for Sentiment
We used the siebert/sentiment-roberta-large-english model from Hugging Face to perform sentence-level sentiment classification on all news articles. For each entity of interest (e.g., Russia, Ukraine, Bucha), we first extracted all sentences from the article that mentioned the entity, using spaCy’s English language model for sentence segmentation.
Prompt:
What is the sentiment(positive, negative, or neutral) towards {entity} in this context: ’{text}’?
A.1.3 Annotation for Article-level Tags
To annotate articles with high-level thematic tags, we used a zero-shot prompt with GPT-3.5 that instructed the model to extract a set of concise, specific tags summarizing each article’s content. Tags were intended to capture relevant people, locations, military units, political themes, and ideological framings. To ensure consistency and support post-processing, the prompt enforced a structured JSON output format.
Prompt:
You are a geopolitical analyst trained to extract concise topical tags from news articles related to international conflict, especially the Russia–Ukraine war. Your task is to read the content of a single article and extract 5 to 12 relevant tags that capture the article’s key topics, actors, places, events, or narrative themes. Each tag should be 1–4 words long, lowercase unless a proper noun, and specific rather than general. Avoid vague tags like “fake”, “news”, or “war” unless central to the article. Tags may include named individuals, military groups, operations, locations, events, or ideological narratives. Output the result in valid JSON using the following structure:
  "article_123": [
    "bucha incident",
    "sergei lavrov",
    "false flag allegations",
    "media manipulation",
    "un security council"
A.1.4 Annotation for Claim Extraction with Summary and Claim Types
We used a structured few-shot prompt with GPT-3.5 to extract and classify claims from news articles. Each prompt instructed the model to identify distinct claims within an article, summarize them neutrally, and assign a claim_type based on a predefined taxonomy. Claims could span one or more sentences and were returned in a structured JSON format to support post-processing.
Prompt:
You are a geopolitical analyst trained to extract claims from news articles about the Russia–Ukraine war. Your task is to read the content of a single article and extract each distinct claim it makes. For each claim, return the original sentence(s), a short neutral summary, and a claim_type. Claims may span multiple sentences. Use only the article content field for analysis. The claim_type must be one of the following: Factual, Attributed, Disputed, Reported, Command/Policy, or Analytical/Assessment. Output the result in valid JSON using the following structure:
  "article_123": {
    "claim_1": {
      "sentence":
        "Ukrainian Side Reports Multiple
        Casualties in Kyiv Airstrike.",
      "summary":
        "Ukraine reports casualties in
        Kyiv airstrike.",
      "claim_type": "Attributed"
    "claim_2": {
      "sentence":
        "Zelensky Orders: ’Ensure NATOs
        Long-Term Presence in the Black
        Sea’.",
      "summary":
        "Zelensky orders long-term NATO
        presence in the Black Sea.",
      "claim_type": "Command/Policy"
A.2 Use Case Experiment Details
Figures 4(a) and 4(b) illustrate 5-day interval favor stance percentages toward Russia and Ukraine across news source countries during March–April 2022.
(a) favor stance towards Russia (by source country).
(b) favor stance towards Ukraine (by source country).
Figure 4: 5-day interval favor stance percentages toward Russia and Ukraine (during the Bucha massacre, March–April 2022). News outlets are grouped by their national affiliations.